OpenStudy (anonymous):

Kevin plans to put concrete on a rectangular portion of his driveway. The portion is 10 feet long and 3 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Kevin needs is $72.59. Which of the following is closest to the width of the portion of the driveway on which Kevin plans to put concrete? [1 foot = 12 inches; 1 yard = 3 feet] 3 feet 6 feet 8 feet 10 feet

OpenStudy (anonymous):

cubic yards= w x 10 x 3 since he pays $72.59, he does have: \[\frac{ $98 }{ yard^3 } = \frac{ $72.59 }{ x }\]which is the same as \[\frac{ yard^3}{$98 } = \frac{ x }{ $72.59 }\]and so we get\[\frac{ yard^3 \times $72.59}{$98 } = x yards\]he has used concrete equalling 72.59/98 = 0.74 cubic yards

OpenStudy (anonymous):

0.75 cubic yards.

OpenStudy (anonymous):

how many feet would it be

OpenStudy (anonymous):

0.75 yard^3 = x ft x 10 ft x 3 inch

OpenStudy (anonymous):

would the answer be 3 feet

OpenStudy (anonymous):

1 yard = 3 feet 1 foot = 12 inches so: 0.75 yard^3 = x yard x 3.333 yard x ___3 inch___

OpenStudy (anonymous):

convert 3 inch t yard plz

OpenStudy (anonymous):

3 inch = 3/12 foot = 0.25 foot 0.25 foot = 0.25/3 yard 0.25 foot = 0.0833 yard .....

OpenStudy (anonymous):

0.75 yard^3 = x yard x 3.333 yard x 0.0833 yard

OpenStudy (anonymous):

0.75 yard^3 = x yard x 0.277 yard^2

OpenStudy (anonymous):

@pony1234 no, because 3 ft are 1 yard and 0.277 x 1yard == 0.277. we need 0.75

OpenStudy (anonymous):

most likely 8 ft

OpenStudy (anonymous):

for 8ft: 2.66 x 0.277 = 0.74 very close to 0.75

OpenStudy (anonymous):

actually i got 0.74 earlier so yes it's 8ft....